A child leaves her book bag on a slide. The bag, which is at the top of the slide, starts from rest and reaches the bottom in 1.72 s. The mass of the book bag is 3.85 kg, the length of the slide is 2.80 m and the angle of incline is 33.5
Answer #1

a) s= ut + 0.5*at^2

u = o m/s

2.8 = 0.5*a*1.72^2

a = 1.893 m/s^2


b) f = mg sin theta - ma = 13.537 N


c) mew = f/(mg cos theta) = 0.43025


d) v = u + a*t = 0 + 1.893*1.72 = 3.256 m/s